OpenStudy (rock_mit182):

Anyone could walk me through , to set up the equation: A flat rectangular piece of aluminum has a perimeter of 62 inches. The lenght is 9 inches longer than the width. Find the width.

OpenStudy (rock_mit182):

@518nad

OpenStudy (518nad):

perimeter = 2L+2W L=9+W 62=2(9+W)+2W 62=18+2W+2W 62-18=4W W=11

OpenStudy (rock_mit182):

omg you are really fast!

OpenStudy (518nad):

it doesnt matter how fast I do it, do you understand?

OpenStudy (rock_mit182):

for sure

OpenStudy (518nad):

okay do another question for me 50 is perimeter Length is 2 times width, what is Length?

OpenStudy (rock_mit182):

perimeter = 2w + 2l l =2w or 2 +w ?

OpenStudy (518nad):

2+w means 2 more than width 2W = w+w , thats 2 times width

OpenStudy (rock_mit182):

anyway: 50 =2w + 2(2w)

OpenStudy (rock_mit182):

right ?

OpenStudy (rock_mit182):

w = 25 /3

OpenStudy (518nad):

but i asked for length

OpenStudy (rock_mit182):

oh sorry

OpenStudy (518nad):

you must read question more carefully, if this was a test, your teacher cant give you any marks for that

OpenStudy (518nad):

even though you did all that work

OpenStudy (518nad):

One of the most important things to solving problems is to completely understand the question. The better you understand the question, the easier it will seem.

OpenStudy (rock_mit182):

ok bro let me see if solved this one right

OpenStudy (518nad):

okay so what is Length

OpenStudy (518nad):

use L=2W

OpenStudy (rock_mit182):

l=25

OpenStudy (518nad):

w=25/3

OpenStudy (518nad):

L=2w =2*(25/3) =?

OpenStudy (jerry45):

nah